If a freshman is selected at random to become student assistant what is the probability of selecting a student with an IQ between 90 and 120


1
Expert's answer
2022-05-04T14:41:52-0400

The office for admission of XYZ University found out that the IQ of incoming 1200 freshmen normally distributed with a mean of 108 and a standard deviation of 10.


"P(90<X<120)=P(X<120)-P(X\le 90)"

"=P(Z<\dfrac{120-108}{10})-P(Z\le\dfrac{90-108}{10})"

"=P(Z<1.2)-P(Z\le-1.8)"

"\approx0.88493-0.03593=0.8490"